OpenStudy (anonymous):

Find the discriminant of the related quadratic function and determine the number of real and imaginary solutions of 2y^2 - 4y + 1 = 0.

OpenStudy (anonymous):

Equation: \[2y ^{2} - 4y + 1 = 0\]

OpenStudy (campbell_st):

the discriminant for a quadratic \[ay^2 + by + c = 0\] can be found using \[\Delta = b^2 - 4ac\] in your question a = 2, b = -4 and c = 1 just substitute and calculate an answer

OpenStudy (campbell_st):

and you may like to ask your teacher, what the question asks how many solutions... and suggest to the teacher a degree 2 polynomial will have 2 solutions...

OpenStudy (anonymous):

@campbell_st - The answer would be: \[(-4)^{2} - 4(2)(1)\] which is equal to 8. If the discriminant is greater than 0, then there are 2 real solutions, right?

OpenStudy (campbell_st):

conditions for the discriminant discriminant > 0 2 real unequal solutions is discriminant is a perfect square the solutions are rational, otherwise the solutions will be real and irrational. discriminant = 0 2 real equal solutions discriminant < 0 2 complex solutions.

OpenStudy (campbell_st):

yes the solutions are unequal and real... do look at 8... is it a square number..? if so the solutions are real, unequal and rational. if not, real unequal and irrational

OpenStudy (anonymous):

@campbell_st - Thank you so much!

OpenStudy (campbell_st):

glad to help